OpenStudy (anonymous):

Simplify: (n + 3)^2 A. n^2 + 9 B. n^2 + 9n + 9 C. n^2 + 6n + 9 D. n^2 + 6

OpenStudy (hoblos):

(n + 3)^2 = n^2 +2(n)(3) + 3^2 = n^2 + 6n + 9

OpenStudy (anonymous):

good answer